OpenStudy (anonymous):

How would I rearrange the following formula so as to solve for a? d=1/2(a)(t^2)+ai(t)

OpenStudy (anonymous):

Would it be a=(d-vi(t))/(1/2(t^2)

OpenStudy (mertsj):

\[d=\frac{1}{2at^2+ait}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

No, \[d= \frac{ 1 }{ 2 }at ^{2}+v _{i}t\]

OpenStudy (mertsj):

1. Subtract v_it from both sides 2. Multiply both sides by 2 3. Divide both sides by t^2

OpenStudy (anonymous):

\[(d-vi)/t^2=\frac{ 1 }{2 }a\]

OpenStudy (anonymous):

So then we get that ^ ?

OpenStudy (anonymous):

An how can we subtract vi if it's attached to t?

OpenStudy (mertsj):

\[d-v _{i}t=\frac{1}{2}at^2\]

OpenStudy (anonymous):

Oh sorry you said vit

OpenStudy (mertsj):

\[2d-2v _{i}t=at^2\]

OpenStudy (mertsj):

\[\frac{2d-2v _{i}t}{t^2}=a\]

OpenStudy (anonymous):

Thank you!!

OpenStudy (mertsj):

yw